A valószínűségszámítás és matematikai statisztika 1. dolgozat (október 28.)

Feladatok

1.  2.  3.  4.  5.  6.
Agócs László Péter  −  +  −  +  +  +
Arany Gábor  −  −  −  −  −  −
Bálint Levente  −  −  +  −  −  −
Csipkés László  ±  −  −  ±  −  −
Eszenyi László  −  +  −  +  −  −
Ferenczik Zoltán  −  +  −  ±  −  −
Fogarasi Edina  −  +  −  −  +  +
Gerják Balázs  ±  +  −  +  +  ±
Juhász Tibor  ±  +  ±  +  +  −
Kis Bertalan  +  +  −  +  +  +
Kocsis Gergely  ±  +  +  ±  +  +
Kovács György  ±  +  +  +  +  −
Lázár Gábor  ±  +  ±  ±  ±  +
Nagy Csaba  −  ±  −  ±  ±  +
Paluska Mihály  −  ±  −  −  ±  −
Pári Viktor  −  ±  −  ±  −  −
Pelejtei Gergely  −  −  −  ±  −  −
Sipos Nándor  −  +  −  −  −  +
Szakácsi János  ±  +  +  +  +  −
Szalontai Gábor  ±  +  ±  +  +  +
Szalontai József  −  +  −  +  −  +
Tóth János  ±  +  +  +  +  +
Tóth Tamás  ±  +  −  +  −  −
Virág Tamás  ±  ±  −  ±  +  −
Zabos Katalin  ±  +  −  ±  −  +

Feladatok

1.  2.  3.  4.  5.  6.
Dancsi József  ±  +  −  −  −  −
Emődi Szabolcs Árpád  ±  ±  −  +±  ±  −
Hamza Péter  −  +  −  +  −  −
Hank Sándor  ±  −  −  ±  +  −
Lörinczi József  ±  +  −  ±  +  +
Molnár Angéla  ±  +  ±  ±  +  +
Molnár Tibor  ±  +  −  +  +  +
Nagy Gábor  ±  +  −  ±  ±  +
Nagy Zsolt  −  −  −  −  −  −
Páli Gábor János  ±  ±  −  −  −  −
Paper Andrea  −  ±  −  ±  ±  +
Papp Gyula  −  ±  −  ±  −  −
Szabó Zsolt  ±  +  ±  ±  −  −
Szevelényi Attila  −  ±  −  ±  −  +
Tilki Csaba  ±  ±  +  ±  +  −
Tordai Jácint  ±  +  −  ±  ±  +